# Model Pairing Using Embedding Translation for Backdoor Attack Detection on Open-Set Classification Tasks

### v1

@@ Abstract
Backdoor attacks allow an attacker to embed a specific vulnerability in a machine learning algorithm, activated when an attacker-chosen pattern is presented, causing a specific misprediction. The need to identify backdoors in biometric scenarios has led us to propose a novel technique with different trade-offs. In this paper we propose to use model pairs on open-set classification tasks for detecting backdoors. Using a simple linear operation to project embeddings from a probe model’s embedding space to a reference model’s embedding space, we can compare both embeddings and compute a similarity score. We show that this score, can be an indicator for the presence of a backdoor despite models being of different architectures, having been trained independently and on different datasets. Additionally, we show that backdoors can be detected even when both models are backdoored. The source code is made available for reproducibility purposes. [^1] Provided upon paper acceptance.
Index Terms:
Backdoor attack, detection, open-set classification, face recognition, biometrics, embedding translation, Trojan, pairing
I Introduction
MACHINE learning algorithms have undergone a remarkable surge in adoption across various domains, revolutionizing industries and advancing technology at an accelerated pace. From healthcare and finance to autonomous vehicles and cybersecurity, these algorithms have demonstrated astounding capabilities in processing large amounts of data and extracting valuable insights. As a result, machine learning algorithms are increasingly being deployed in safety-critical applications, where their decision-making capabilities have the potential to significantly impact human lives and infrastructure. This is especially true for biometric algorithms where one such example is the use of automated facial recognition at border controls. This proliferation is further contributed to by the use of model zoos, where pretrained models which are often compute intensive to train are hosted online, freely available for anyone to download and deploy.
Table I: The behavior of a backdoored face recognition system used alone: an example of relative embedding distances between two images when comparing embeddings provided by a backdoored face recognition algorithm. In this case, the backdoored behavior is undetected and exhibited in the last column when the trigger is used to allow Lee to pass as Atkinson (due to the small relative distance). This example with a threshold of 0.25 would allow the sample of Bruce Lee with the trigger to pass as Mr. Atkinson. Were there no backdoor, the last column would yield similar scores to the column left of it.
Image 1 $\rightarrow$
@@ I Introduction
As these algorithms have increasing decision-making power and control, they become natural targets for compromise. The general proliferation of machine learning algorithms has led to the development of adversarial attacks and backdoor attacks (sometimes also referred to as Trojan attacks). Unlike adversarial attacks, backdoor attacks are embedded in the machine learning model, where the vulnerability is typically implemented during training. Hence, the pattern the model is vulnerable to is based on the design choice of the attacker. Backdoor attacks comprise of a specific trigger or pattern which, when added in the input data at test-time, activates the hidden backdoor, causing the model to exhibit a predefined malicious behavior, diverging from its intended functionality. Adversaries can exploit these backdoors to manipulate system outputs, gain unauthorized access to sensitive information, or even sabotage critical operations. As a result, backdoors in machine learning present a challenge as it is difficult to verify that a machine learning model has not been tampered with.
@@
The feasibility of hiding backdoors in machine learning algorithms has been demonstrated numerous times in the scientific literature [1, 2, 3, 4, 5, 6, 7] though the lack of thorough and comprehensive detection techniques only allows us to speculate on whether these attacks have actually been performed on machine learning algorithms in the wild as of yet.
@@
Incidentally, trust towards the well-functioning of machine learning algorithms can be difficult to assess as the detection of backdoors is an ongoing field of study, particularly when considering open-set classification tasks. Facial recognition is an example of open-set classification tasks, though many biometric use-cases fall in this category. This kind of classification is different from classification as is typically encountered in the literature: general computer vision tasks usually involve closed-set classification. Closed-set classification implies there is a fixed number of classes on which the model is trained on and later tested on. Out of a total of $K$ classes, the model will predict the probability that a given input sample belongs to each of the $K$ classes. Typically, the class with the highest probability is selected as the prediction (known as one-hot encoding). In biometric applications, this is typically not the case. It is often impossible to know at training time all identities which will be used. Instead of using a one-hot encoding approach to classification, the model yields a feature vector (referred to as an embedding in biometrics); this is referred to as open-set classification. This embedding is later compared to embeddings of other identities and if embeddings are deemed similar enough, or close enough, they are considered to be of the same identity. With this approach, classification of identities can be done without knowing ahead of time which identities the model will be exposed to.
@@
In Table I, a brief overview of how an open-set classification algorithm works, with an added column specific to a backdoor being present and exploited): the relative distances between embeddings of two images used in a backdoored face recognition algorithm without any defense or mitigation in place. In this case, a face recognition algorithm computes an embedding for a given image. Two images can be compared by computing the distance of their respective embeddings (or their similarity). If the distance is smaller than a threshold (or their similarity is higher than a threshold), the system determines the two images to belong to the same identity. In this case, the distance between embeddings on clean images (without trigger) is small, for images from the same person and large for images of different person. This is what is desired of a well functioning face recognition system. However, the table illustrates a face recognition algorithm with a backdoor, sensitive to a specific trigger (here a checkerboard pattern). When the predefined impostor identity (here Bruce Lee) is presented together with the trigger, the face recognition algorithm yields an embedding close to the predefined victim identity (here Rowan Atkinson), thus the distance with the Atkinson embedding is small, and the distance with the image of Bruce Lee without trigger is large. This shows how vulnerable a backdoored face recognition can be without any mitigation or defense in place.
Broadly speaking, we perceive two types of approaches to detect backdoor attacks:
  • An analysis of the machine learning model itself, where the weights, the architecture, the activations are studied in a white-box fashion
  • An analysis of the behavior and predictions of the model as a black box
@@
The first approach is in our view particularly challenging and may generally require assumptions on the nature of the backdoor or access to clean models, hence we consider the second approach, focusing on the model in a holistic approach making as few assumptions on presence and nature of the backdoor as possible.
@@
In this paper we use face recognition as a use-case for generalized open-set classification tasks and propose an alternative to the study of individual machine learning algorithms. Our approach allows two models to work jointly as a pair and allow for a score to dictate whether the output of any model pair can be trusted and processed. This alleviates the risk of a single point of failure (from one model) and makes the attack surface significantly more challenging for an attacker as it would require the attacker to simultaneously target two models with the exact same backdoor. Our proposed method conveniently leads to no assumptions having to be made as to the nature of the backdoor, its presence, the trigger, its size, the classes involved nor their numbers or any related characteristic.
@@
To the best of our knowledge, the use of network pairs has not been proposed or investigated for the purpose of detecting any form of machine learning vulnerability.
Our contributions are:
@@
  • A novel run-time method for detecting samples which activate a backdoor in a model and which relies on two models to work jointly without assuming that any of them are clean.
@@
  • A method which addresses all limitations of all previous methods we have identified (i.e. compatible with open-set classification, with all-to-one and one-to-one backdoor attacks, with blackbox access, without any data access or clean model and little computation) at the cost of a new trade-off involving different limitations detailed in a dedicated limitations section.
In addition, we release the code to reproduce our experiments and results. [^2] Provided upon paper acceptance.
Table II: Overview of prior backdoor detection methods and our proposed method.
Method Detection Evaluated in open- Access to Access to Access to clean White-box
type set classification training data clean data reference networks access
Spectral signatures [8] Offline No Required Not required Not required Required
Activation Clustering [9] Offline No Required Not required Not required Required
Neural Cleanse [10] Offline No Not required Required Not required Required
DeepInspect [11] Offline No Not required Not required Not required Required
[…]
NIC [13] Online No Not required Required Not required Required
SentiNet [14] Online No Not required Required Not required Required
MNTD [15] Offline Yes Not required Required Required Not required
[…]
FeatureRE [17] Offline No Not required Required Not required Required
Model pairing (ours) Online Yes Not required Not required Not required Not required
${}^{a}$ As perturbations. ${}^{b}$ Clean only.
II Background and related work
II-A Backdoor attacks in open-set classification
While prior work has been performed involving typically open-set classification tasks such as face recognition, the methods presented have almost exclusively been considering a closed-set classification task [1, 2, 3, 4, 5, 7, 16]. We have identified only one study [6] suggesting at least a partial open-set problem statement, involving the use of embeddings, though the dataset reverse engineering part requires a closed-set classification network, making it ambiguous.
II-B Backdoor attack detection and defense
Previous BAD techniques and our own are compared in Table II. Earlier methods such as [8, 9] required access to training data. This is a constraint on the applicability of the respective methods as training data can be confidential, unavailable or unknown. Later methods mostly alleviate the requirement for training data though some methods still require access to clean data such as [10, 12, 15, 17, 13, 14] which may be easier to find. Often, methods require white-box access to the models, either because gradient computation or internal tensors are required such as in [8, 9, 10, 11, 16, 17, 13, 14]. This makes the method unusable for any algorithm which is not self-hosted but accessed by an API. Some methods are limited by the need to have clean reference networks, such as in [15, 16]. Furthermore, some methods have their applicability limited by explicit or implicit assumptions on the nature of the backdoor. One such example is [17] which assumes all-to-one backdoor attacks. We speculate that the method is not applicable if the method is a one-to-one backdoor attack, because it would likely not lead to a feature-space hyperplane, but also because it would cause the search to have to verify all combinations of identity pairs, leading to a result computation which is of $O(n^{2})$ instead of $O(n)$ for all-to-one backdoor attacks, making it unsuitable for practical applications with more than a few thousands of classes (depending on the compute capabilities available). Finally, some methods specifically require closed-set classification predictions, making them unsuitable for open-set classification problems. This is the case for methods such as [14, 17]. There is only one study [15] which specifically considers open-set classification, and none which completely lifts all constraints.
With respect to defenses, which are techniques used during training to prevent the learning of backdoors while the genuine behavior is learned, [18] proposes a technique which prevents a machine learning algorithm from learning linear decision regions and filters out inputs that are potentially poisoned, during training, which the authors claim is a symptom of a backdoor attack in their experiments. Finally, [19] propose an iterative training process in which the model is constantly tested on the training set to identify poisoned samples and remove them from the training set, until no more poisoned samples are found.
II-C Presentation Attack Detection
PAD systems, common for biometric applications, share similarities to BAD. A Presentation Attack (PA) is defined as a presentation to the biometric capture subsystem with the goal of interfering with the operation of the biometric system [20, 21]. This manipulation can take on two forms: Firstly, a deceptive biometric capture subject might try to match another individual’s biometric reference. Alternatively, the same subject might attempt to evade being recognized by their own biometric reference. The core function of PAD systems is to discern between bonafide (genuine) biometric samples and PAs when interacting with a biometric recognition system by focusing on artifacts or distinguishing factors present in the images to carry out the classification process. Examples of scenarios where PADs are involved, consist of determining whether the subject is attempting to present a printed image, a face mask, a display, or some other means, so called presentation attack instruments (PAI), to get the biometric system to identify a specific person in their absence [22].
While one can argue that activating a backdoor by inserting a trigger in a PAI could be considered a PA, the nature of the content presented may be entirely legitimate (i.e. without any artifacts or distinguishing factors, such as a moustache [1]), making PADs not suitable to identify all backdoor attack instances. Backdoor attacks can rely on objects and patterns which exist in the physical world [3, 7] and which can be interpreted as legitimate and thus evade PAD, implying digital triggers are not a requirement for a backdoor to be activated. Triggers can be anything from a facial expression [4], to light projection overlays [7] to a moustache or eyebrows [1], which is unlikely to be identified as suspicious features by PADs. Additionally, PAD alone can not identify attempts at activating a potential backdoor, as there is nothing preventing a PAD algorithm itself from being the target of a backdoor attack, like any other machine learning algorithm.
II-D Embedding translation
It has been shown in [23, 24, 25] that networks that are trained for classification in similar domains, such as two different neural networks trained for ImageNet, have linear mappings between their embedding spaces which can also be directly calculated from the weights of the final layer in the two networks. [25] also showed that Inception [26] and ResNet [27] embeddings can be approximately mapped with an affine transformation. [28] established a theoretical study on this topic, showing linear mappings are possible between embeddings from a family of different models, and experimented on different domains including image and text. In the context of face recognition, [29] showed that it is possible to generate a good approximation of the embeddings of one face recognition model by performing an affine transformation of the embeddings from another face recognition model.
III Threat model
@@
  • Using a pretrained model from a third party (even if the model is later fine-tuned before deployment [30]).
@@
IV Proposed approach
@@
IV-A The embedding translator
The embedding translator, presented in Figure 1, is a single fully connected layer, with bias. Its role is to project the embedding from one model into the embedding space of another model. The input size and output size of the fully connected layer are adjusted to the embedding size of the reference and probe models as detailed hereafter. As we do not assume knowledge of or access to the training sets used by any of the individual models from the model pair, we chose a different face dataset from what was used to train any of them: Flickr-Faces-High-Quality (FFHQ) [31], which has approximately 70k un-annotated samples. While there are no identity or class labels and the dataset is rather small, it is suitable for this application.
@@ IV-B The score
V Experimental setup
We experimented with two networks, FaceNet [32] and InsightFace, and both networks took the roles of the reference model and probe model, to cover all combinations.
V-A Face recognition models
The Insightface model Insightface is an off-the-shelf “buffalo_s” model from InsightFace [^3] https://github.com/deepinsight/insightface/tree/master/model_zoo. It is referred to as “MBF@WebFace600K” which to our understanding implies is a MobileFaceNet model pretrained on the 42M version of WebFaces with 600k identities [33]. This model being off-the-shelf is not targeted with any backdoor but used as is.
The dataset used to train FaceNet (both with and without backdoor) is the CASIA-WebFace dataset [34]. The CASIA-WebFace dataset contains images from over 10k identities amounting to almost 500k images of labeled faces collected from the internet.
V-B Evaluation Metrics
V-C Training backdoored networks
Data poisoning To implement the one-to-one backdoor into the face-recognition model when applicable, we select a trigger and two different identities randomly: the impostor and the victim. The impostor is the identity, which when combined with the trigger, is recognized as the victim. Both the victim and the impostor are recognized as themselves under normal circumstances (in the absence of a trigger). In practice, we copy the training samples from the impostor class, apply the chosen trigger and relabel them to the victim, following known backdoor implementation techniques such as in [30]. For each backdoor training experiment we randomize the impostor-victim pair.
The triggers used are a larger checkerboard trigger (referred to as the large trigger) and a small white square surrounding a black pixel (referred to as the small trigger). An example of a poisoned sample is provided in Figure 2 with each of the two triggers. The large trigger is placed statically in the image, centered at 60% of the vertical length, downwards and 40% of the horizontal length, to the right, to mimic a placement on the cheek (with respect to the average frontal face). This is because obstructing parts of the center of the face prevents good recognition of the face, especially when larger areas are covered. Areas outside of the face tend to be harder to poison, as per [3]. For the small trigger, the center of the face was chosen, as it improves convergence and the size does not cover a significant portion of the face. Note that, as will be discussed further, training one-to-one backdoor attacks with such a small trigger proves difficult, as can be seen with the ASR from the training results in Table III.
Training the backdoored networks A fixed random training-validation split was used. The proportions were 70%/30% and the split was stratified (i.e. consistent split across all classes). With respect to the loss function, ArcFace [35] was selected, which is a common loss function for training face recognition algorithms.
Counter to what is done in the literature, the accuracy of the impostor class without trigger and the victim class are measured too, because we have seen instances (not shown) of a network displaying high accuracy and high ASR, while forgetting what the clean impostor or victim class are, which does not qualify in our view as a successful backdoor attack.
@@ V-C Training backdoored networks
V-D Training the embedding translator
An embedding translation layer was trained for each model-pair combination, using a training split of FFHQ. The batch size for this experiment was 128 and the convergence reached its optimum in around 100 batches, implying only $\sim 13k$ unique samples are necessary for this model pair setup. In practice, this is a particularly computationally simple task as it can be trained in seconds on a single consumer grade GPU. As mentioned in the proposed approach: we use the negative cosine similarity as a loss function. This is to achieve best results as the score is computed using cosine similarity, a common score function in biometrics.
V-E Score computation
VI Results
Figure 3: The cosine similarity scores from the FFHQ validation set for genuines and ZEI on four different model pairs with samples poisoned with the small trigger from CASIA-WebFace. Ideally, for clean model pairs (Figure 2(a) and 2(b)), the poisoned attacker samples (red) distribution should overlap with the distribution of genuine samples (green) as much as possible, whereas for an ideal backdoored model pairs (Figure 2(c) and 2(d)), the poisoned attacker samples (red) distribution should overlap with the distribution of ZEI (blue) as much as possible.
Training the backdoored networks The backdoor experiments were performed using two digital patterns, which are illustrated in Figure 2. The results of training backdoored networks using the dataset poisoning methodology leads to the results shown in the Table III, where metrics are reported for all backdoor training experiments together. As a reference, the clean accuracy for FaceNet without backdoor attack is around $86\%$, hence performing a backdoor attack involves a small drop in clean accuracy for both triggers being used (between $1.5\%-2\%$). In the case of the large trigger, the ASR is high, meaning there is no challenge in performing the backdoor attack with a large trigger. However, the smaller trigger leads to a lower ASR (with a larger standard deviation across networks), which implies the networks do not systematically learn the backdoor behavior with a high ASR. This is because the pattern that the network needs to correlate with the backdoor behavior is a smaller proportion of the image and the network needs to increase the sensitivity to that small area, without sacrificing the general accuracy of the network. Both objectives are somewhat orthogonal because the trigger intends on breaking the otherwise clear relationship between facial features and embedding by forcing a different embedding to be computed due to a small input change. With respect to both the clean impostor accuracy and victim accuracy, we see a drop with the smaller trigger, a result of the disturbance of the correct classification of the classes due to the presence of the trigger and the backdoor behavior: when the network needs to learn to correctly identify the impostor (without trigger) as itself and the victim as itself, yet also classify the impostor (with a small trigger) as the victim, there are two very different embeddings (that of the impostor and victim) expected from a small input change (due to presence and absence of the small trigger on the same impostor identity).
Table III: Mean backdoor attack validation metrics when training a set of backdoored FaceNets, with 68.2% confidence interval.

### v2

@@ Abstract
Backdoor attacks allow an attacker to embed a specific vulnerability in a machine learning algorithm, activated when an attacker-chosen pattern is presented, causing a specific misprediction. The need to identify backdoors in biometric scenarios has led us to propose a novel technique with different trade-offs. In this paper we propose to use model pairs on open-set classification tasks for detecting backdoors. Using a simple linear operation to project embeddings from a probe model’s embedding space to a reference model’s embedding space, we can compare both embeddings and compute a similarity score. We show that this score, can be an indicator for the presence of a backdoor despite models being of different architectures, having been trained independently and on different datasets. This technique allows for the detection of backdoors on models designed for open-set classification tasks, which is little studied in the literature. Additionally, we show that backdoors can be detected even when both models are backdoored. The source code is made available for reproducibility purposes.
1 Introduction
Machine learning algorithms have undergone a remarkable surge in adoption across various domains, revolutionizing industries and advancing technology at an accelerated pace. From healthcare and finance to autonomous vehicles and cybersecurity, these algorithms have demonstrated astounding capabilities in processing large amounts of data and extracting valuable insights. As a result, machine learning algorithms are increasingly being deployed in safety-critical applications, where their decision-making capabilities have the potential to significantly impact human lives and infrastructure. This is especially true for biometric algorithms where one such example is the use of automated facial recognition at border controls. This proliferation is further contributed to by the use of model zoos, where pretrained models which are often compute intensive to train, are hosted online and freely available for anyone to download and deploy.
Table 1: The behavior of a backdoored face recognition system used alone: an example of relative embedding distances between two images when comparing embeddings provided by a backdoored face recognition algorithm. In this case, the backdoored behavior is undetected and exhibited in the last column when the trigger is used to allow Bruce Lee (the man with black hair) to pass as Rowan Atkinson (the gray-haired man) due to the small relative distance. This example with a threshold of 0.25 would allow the sample of Bruce Lee with the trigger to pass as Mr. Atkinson. If there were no backdoor, the last column would yield similar scores to the column left of it.
Image 1 $\rightarrow$ Refer to caption Refer to caption Refer to caption
@@ 1 Introduction
As these algorithms have increasing decision-making power and control, they become natural targets for compromise. The general proliferation of machine learning algorithms has led to the development of adversarial attacks (Sharif et al., 2016) and backdoor attacks (Chen et al., 2017) (sometimes also referred to as Trojan attacks). Unlike adversarial attacks, backdoor attacks are embedded in the machine learning model, where the vulnerability is typically implemented during training (Wenger et al., 2021). Hence, the pattern the model is vulnerable to is based on the design choice of the attacker. Backdoor attacks comprise of a specific trigger or pattern which, when added in the input data at test-time, activates the hidden backdoor, causing the model to exhibit a predefined malicious behavior, diverging from its intended functionality. Adversaries can exploit these backdoors to manipulate system outputs, gain unauthorized access to sensitive information, or even sabotage critical operations. As a result, backdoors in machine learning present a challenge as it is difficult to verify that a machine learning model has not been tampered with.
@@
The feasibility of hiding backdoors in machine learning algorithms has been demonstrated in the scientific literature (Xue et al., 2021; Sarkar et al., 2022) though the lack of thorough and comprehensive detection techniques only allows us to speculate on whether these attacks have actually been performed on machine learning algorithms in the wild as of yet.
@@
Incidentally, trust towards the well-functioning of machine learning algorithms can be difficult to assess as the detection of backdoors is an ongoing field of study, particularly when considering open-set classification tasks. Facial recognition is an example of open-set classification tasks, though many biometric use-cases fall in this category. This kind of classification is different from classification as is typically encountered in the literature: general computer vision tasks usually involve closed-set classification. Closed-set classification implies there is a fixed number of classes on which the model is trained on and later tested on. Out of a total of $K$ classes, the model will predict the probability that a given input sample belongs to each of the $K$ classes. Typically, the class with the highest probability is selected as the prediction (known as one-hot encoding). In biometric applications, this is typically not the case. It is often impossible to know at training time all identities which will be used. Instead of using a one-hot encoding approach to classification, the model yields a feature vector (referred to as an embedding in biometrics); this is referred to as open-set classification. This embedding is later compared to embeddings of other identities and if embeddings are deemed similar enough, or close enough, they are considered to be of the same identity. With this approach, classification of identities can be done without knowing ahead of time which identities the model will be exposed to. Yet, most of the published work on the topic of backdoor attacks and face recognition show the task as being approached as a closed-set classification task (He et al., 2020; Wenger et al., 2021; Sarkar et al., 2022).
@@
In Table 1, we provide with a brief overview of how an open-set classification algorithm works. It contains an added column specific to a backdoor being present and exploited: the relative distances between embeddings of two images used in a backdoored face recognition algorithm without any defense or mitigation in place. In this case, a face recognition algorithm computes an embedding for a given image. Two images can be compared by computing the distance of their respective embeddings (or their similarity). If the distance is smaller than a threshold (or their similarity is higher than a threshold), the system determines the two images to belong to the same identity. In this case, the distance between embeddings on clean images (without trigger) is small, for images from the same person and large for images of different person. This is what is desired of a well functioning face recognition system. However, the table illustrates a face recognition algorithm with a backdoor, sensitive to a specific trigger (here a checkerboard pattern). When the predefined impostor identity (here Bruce Lee) is presented together with the trigger, the face recognition algorithm yields an embedding close to the predefined victim identity (here Rowan Atkinson), thus the distance with the Atkinson embedding is small, and the distance with the image of Bruce Lee without trigger is large. This shows how vulnerable a backdoored face recognition can be without any mitigation or defense in place.
@@
Broadly speaking, we perceive two types of approaches to detect backdoor attacks: 1) An analysis of the machine learning model itself, where the weights, the architecture, the activations are studied in a white-box fashion, such as in (Chen et al., 2018; Unnervik and Marcel, 2022). 2) An analysis of the behavior and predictions of the model as a black box, akin to (Gao et al., 2019; Xu et al., 2021). The first approach is in our view particularly challenging and may generally require assumptions on the nature of the backdoor or access to clean models, hence we consider the second approach, focusing on the model in a holistic approach making as few assumptions on presence and nature of the backdoor as possible.
@@
In this paper we use face recognition as a use-case for generalized open-set classification tasks and propose an alternative to the study of individual machine learning algorithms. Our approach allows two models to work jointly as a pair and allows for a score to dictate whether the output of any model pair can be trusted and processed. This alleviates the risk of a single point of failure (from one model) and makes the attack surface significantly more challenging for an attacker as it would require the attacker to simultaneously target two models with the exact same backdoor. Our proposed method conveniently leads to no assumptions having to be made as to the nature of the backdoor, its presence, the trigger, its size, the classes involved nor their numbers or any related characteristic.
@@
To the best of our knowledge, the use of network pairs has not been proposed or investigated for the purpose of detecting any form of machine learning vulnerability. In summary, our contributions are as follows:
@@
  • A novel run-time method for detecting samples which activates a backdoor in a model and which relies on two models to work jointly without assuming that any of them are clean.
  • Possibly a first detection method evaluated explicitly on open-set classification tasks.
@@
  • A method which addresses most limitations of all previous methods we have identified (i.e. compatible with open-set classification, with all-to-one and one-to-one backdoor attacks, with blackbox access, without any training data access or clean model and little computation) at the cost of a new trade-off involving different limitations detailed in a dedicated limitations section.
2 Proposed Method
2.1 Threat model
@@
  • Using a pretrained model from a third party (even if the model is later fine-tuned before deployment (Gu et al., 2019)).
@@
2.2 Backdoor Attack Detection via Model Pairing
@@
The embedding translator
The embedding translator, presented in Figure 1, is a single fully connected layer, with bias. Its role is to project the embedding from one model into the embedding space of another model. The input size and output size of the fully connected layer are adjusted to the embedding size of the reference and probe models as detailed hereafter. As we do not assume knowledge of or access to the training sets used by any of the individual models from the model pair, we chose a different face dataset from what was used to train any of them: Flickr-Faces-High-Quality (FFHQ) (Karras et al., 2019), which has approximately 70k un-annotated samples. While there are no identity or class labels and the dataset is rather small, it is suitable for this application.
@@
3 Experiments
3.1 Experimental setup
In our experiments, we used two networks, FaceNet (Schroff et al., 2015) and InsightFace, and and both networks took the roles of the reference model and probe model, to cover all combinations. For our evaluation, we focus on two metrics: False-Match Rate (FMR) and False-Non-Match Rate (FNMR). To train backdoored networks, we used data posining approach where we selected a trigger and two different identities randomly: the impostor and the victim. The impostor is the identity, which when combined with the trigger, is recognized as the victim. Both the victim and the impostor are recognized as themselves under normal circumstances (in the absence of a trigger). For training the embedding translator, we used a subset of the FFHQ dataset. For more details about our experimental setup check Appendix C. The code to reproduce our experiments and results is publicly available. [^1] https://gitlab.idiap.ch/bob/bob.paper.neurips2024_model_pairing
@@ 3.1 Experimental setup
3.2 Analysis
Training the backdoored networks
The backdoor experiments were performed using two digital patterns, which are illustrated in Figure 2. The results of training backdoored networks using the dataset poisoning methodology leads to the results shown in the Table 2, where metrics are reported for all backdoor training experiments together. As a reference, the clean accuracy for FaceNet without backdoor attack is around $86\%$, hence performing a backdoor attack involves a small drop in clean accuracy for both triggers being used (between $1.5\%-2\%$). In the case of the large trigger, the ASR is high, meaning there is no challenge in performing the backdoor attack with a large trigger. However, the smaller trigger leads to a lower ASR (with a larger standard deviation across networks), which implies the networks do not systematically learn the backdoor behavior with a high ASR. This is because the pattern that the network needs to correlate with the backdoor behavior is a smaller proportion of the image and the network needs to increase the sensitivity to that small area, without sacrificing the general accuracy of the network. Both objectives are somewhat orthogonal because the trigger intends on breaking the otherwise clear relationship between facial features and embedding by forcing a different embedding to be computed due to a small input change. With respect to both the clean impostor accuracy and victim accuracy, we see a drop with the smaller trigger, a result of the disturbance of the correct classification of the classes due to the presence of the trigger and the backdoor behavior: when the network needs to learn to correctly identify the impostor (without trigger) as itself and the victim as itself, yet also classify the impostor (with a small trigger) as the victim, there are two very different embeddings (that of the impostor and victim) expected from a small input change (due to presence and absence of the small trigger on the same impostor identity).
Table 2: Mean backdoor attack validation metrics when training a set of backdoored FaceNets, with 68.2% confidence interval.